if % abundance of two isotopes C-12 and C-14 are 90% and 10% respectively, then number of C-12 atoms in 12 grams of sample will be?

Grade:11

1 Answers

Beenu Mathew
7 years ago
FIRST YOU HAVE TO CALCULATE THE ATOMIC MASS USING THE RELATIVE ABUNDANCES GIVEN TO YOU
HOW TO DO THAT
ATOMIC MASS = 90/100 * 12 + 10/100 * 14
                        =12.2g
now,given mass is 12g
so 12/12.2g gives you the number of moles
12/12.2 * Na(avogardo number) = 5.9 * 1023 (appx)
PLS REVERT BACK IF CORRECT,
ALSO IF CORRECT PLS APPROVE
THANK YOU,
BEST OF LUCK
THANK YOU